OpenStudy (anonymous):

Find x and v in terms of c and b. (x/c)+(v/b)=1 (x/b)+(v/c)=1

OpenStudy (asnaseer):

first multiply each equation by bc to get rid of the fractions. so 1st equation simplifies to:\[bc*(\frac{x}{c}+\frac{v}{b})=bc*1\]\[bx+cv=bc\]and second simplifies to:\[bc*(\frac{x}{b}+\frac{v}{c})=bc*1\]\[cx+bv=bc\]so now you ave two equations in two unknowns (x & v):\[bx+cv=bc\]\[cx+bv=bc\]solve using elimination method.